# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### v1

@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
Finally, we estimate $J_{8}$. Recalling the notation (3.9), we denote by
$V_{w,i}\coloneqq U_{w}-u_{i},\qquad i=1,2,$ (3.37)
and hence $V_{w,2}-V_{w,1}=\delta u$. Thus, by using identity (3.11), one can compute, for $i=1,2$,
@@
$\|\partial_{i}(\mathcal{T}_{w,1}-\mathcal{T}_{w,2})\|_{L^{2}}\lesssim(\|V_{w,1}\|_{L^{\infty}}+\|V_{w,2}\|_{L^{\infty}})\|\partial_{i}\delta u\|_{L^{2}}+(\|\partial_{i}V_{w,1}\|_{L^{4}}+\|\partial_{i}V_{w,2}\|_{L^{4}})\|\delta u\|_{L^{4}}.$ (3.38)
Hence one has by using (3.37)
$|J_{8}|$ $\lesssim\lVert\nabla\delta u\rVert_{L^{2}}\bigg{[}\big{(}\lVert U_{w}-u_{1}\rVert_{L^{\infty}}+\lVert U_{w}-u_{2}\rVert_{L^{\infty}}\big{)}\lVert\nabla\delta u\rVert_{L^{2}}$ (3.39)
@@
Thus, we have derived the following from estimate (3.36),
$\begin{gathered}\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg{]}\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}).\end{gathered}$ (3.40)
Consequently, combining estimates (3.35) and (3.40) leads to
$\begin{gathered}\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert\delta u\rVert_{L^{2}}^{2}+\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}(1+\alpha^{2})\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg{]}\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}),\end{gathered}$ (3.41)
from which the stability (i.e. the continuous dependence on the initial data) and uniqueness of the strong solution follows. This concludes the proof of the global well-posedness of the intermediate model (1.13).
@@
Thanks to (3.32), where the estimate is independent of $\epsilon$, one can pass to the limit $\epsilon\rightarrow 0$ for a subsequence of solutions to the intermediate system (1.13), denoted as $\{(u_{\epsilon},\sigma_{\epsilon})\}$, with the convergence in the same spaces as in equations (3.26)–(3.31) with $N\rightarrow\infty$ replaced by $\epsilon\rightarrow 0$. The limit is denoted as $(u,\sigma)$, and the limit satisfies estimate (3.32). It is left to verify that $(u,\sigma)$ is a solution to the Voigt-EVP model (1.12).
@@
where we have used Lemma 2.3, the convergence $u_{\epsilon}\rightarrow u$ in $C([0,T];H^{1}(\mathbb{T}^{2}))$ and the elementary inequality $\sqrt{y+z}\leq\sqrt{y}+\sqrt{z}$ for $y,z\geq 0$. It is clear that the linear terms in equations (1.13a)-(1.13b) converge. We now show that the nonlinear terms in the equation for the stress tensor converge. In particular, for all $t\in[0,T]$ we have
@@
In a similar fashion, one can show that the terms $-\frac{3}{2}\frac{\sqrt{\lvert D(u_{\epsilon})\rvert^{2}+\epsilon^{2}}}{P}\operatorname{Tr}\sigma_{\epsilon}\mathbb{I}_{2}$ and $\frac{\sqrt{\lvert D(u_{\epsilon})\rvert^{2}+\epsilon^{2}}}{2}\mathbb{I}_{2}$ converge to $-\frac{3}{2}\frac{\lvert D(u)\rvert}{P}\operatorname{Tr}\sigma\mathbb{I}_{2}$ and $\frac{\lvert D(u)\rvert}{2}\mathbb{I}_{2}$ in $L^{\infty}((0,T);L^{2}(\mathbb{T}^{2}))$. Finally, we show that the ocean stress term $\mathcal{T}_{w}^{\epsilon}$ converges. We find
@@
[…]
@@
To show the (global) well-posedness, one can repeat similar arguments as in Section 3.3. Consequently, one can obtain the analogous version of estimate (3.41), and therefore conclude the stability and uniqueness of the Voigt-EVP model. This finishes the proof of Theorem 1.2.
@@
We remark that the proof of the convergence of the solutions in the viscosity cutoff limit $\epsilon\rightarrow 0$, is in some sense related to the approach in [3], in which the local well-posedness of a model for non-Newtonian incompressible flows with power-law stress tensor was established uniformly in the cutoff parameter of the viscosity.
@@ 5 Discussion and conclusion
In this paper we commenced the rigorous analytical study of one of the fundamental models of sea-ice dynamics, namely the elastic-viscous-plastic (EVP) sea-ice model. We introduced a regularised version of the EVP model, by means of the Voigt regularisation, which is inspired by the original elastic regularisation of the Hibler model in [35] (as both regularisations formally do not impact the asymptotic dynamics). We have shown this Voigt-EVP model to be globally well-posed. One interesting question, which we leave to future work, is whether the Voigt-EVP model also has a suitable notion of weak solution.
Due to the (formal) linear ill-posedness of the EVP model as discussed in Section 1.2, it seems necessary to introduce a regularised version of the EVP model of the type (1.12a)-(1.12c), in order to obtain a well-posed sea-ice model. Finally, in order to conclude the proof of the global well-posedness we took the limit of the viscosity cutoff parameter $\epsilon$ going to zero (while keeping the Voigt regularisation), and obtain that the Voigt-EVP model is globally well-posed.
Moreover, we would like to mention that over time several related models to the EVP model have been developed in the sea-ice literature, see for example [34, 7]. The approach developed here will also have potential applications in the rigorous study of several of these related sea-ice models.
@@
If one includes an advection term in the momentum equation of the intermediate system (1.13a)-(1.13c), by applying the same methods as in this paper, one can prove the local well-posedness of the intermediate system by using similar estimates as in this paper. Note however that the existence time will most likely depend on $\epsilon$ (unlike for Corollary 1.4). This therefore prohibits taking the limit $\epsilon\rightarrow 0$, as we have done in the proof of Theorem 1.2.
@@
We remark that the replacement of $\overline{\mathcal{D}}$ by $\lvert D(u)\rvert$ in the Voigt-EVP model is done merely for the sake of convenience. In fact, by using the approach from this paper one can prove the global well-posedness of the model (1.13a)-(1.13c) in which $\mathcal{D}_{\epsilon}$ has been replaced by
@@
$\left\lVert\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right\rVert_{L^{2}}\lesssim\lVert\nabla u\rVert_{L^{2}}+\epsilon,\quad\left\lVert\nabla\left(\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right)\right\rVert_{L^{2}}\lesssim\lVert\Delta u\rVert_{L^{2}}.$
@@
We also note that, in addition to (1.14), different regularisations of the strain rate have been used in the sea-ice literature. For example, one can set maximal values for the shear and bulk viscosities, as is done in [32]. This means that, instead of $\mathcal{D}_{\epsilon}$, one uses
$\max\{\overline{\mathcal{D}},\epsilon\}=\frac{\overline{\mathcal{D}}+\epsilon}{2}+\frac{1}{2}\lvert\overline{\mathcal{D}}-\epsilon\rvert.$ (5.1)
@@
$\overline{\mathcal{D}}_{\epsilon,\gamma}\coloneqq\frac{\overline{\mathcal{D}}+\epsilon+\gamma}{2}+\frac{1}{2}\sqrt{\lvert\overline{\mathcal{D}}-\epsilon\rvert^{2}+\gamma^{2}}.$ (5.2)
@@
It is straightforward to repeat the proof of Theorem 1.2 to obtain a global strong solution of the Voigt-EVP model with the viscosity regularisation from equation (5.1) by taking the limit $\gamma\rightarrow 0$. Moreover, one can also choose to regularise the viscosities by means of the hyperbolic tangent, as was introduced in [48], see also [17], but again one can treat this regularisation by similar methods as in this paper. Generally speaking, one can treat any $C^{2}$ approximation of $\overline{\mathcal{D}}$ by relying on the approach of this paper.

### v2

@@ 3.3 Well-posedness of the intermediate system (13)
Finally, we estimate $J_{8}$. Recalling the notation (33), we denote by
$V_{w,i}\coloneqq U_{w}-u_{i},\qquad i=1,2,$ (64)
and hence $V_{w,2}-V_{w,1}=\delta u$. Thus, by using identity (35), one can compute, for $i=1,2$,
@@
Note that in the above computation we have used the following relation
$\frac{V_{w,1}\cdot\partial_{i}V_{w,1}}{|V_{w,1}|}V_{w,1}$ $-\frac{V_{w,2}\cdot\partial_{i}V_{w,2}}{|V_{w,2}|}V_{w,2}=\frac{V_{w,1}\cdot\partial_{i}V_{w,1}}{|V_{w,1}|}(V_{w,1}-V_{w,2})+\frac{V_{w,1}\cdot\partial_{i}(V_{w,1}-V_{w,2})}{|V_{w,1}|}V_{w,2}$
$+\bigg(\frac{1}{|V_{w,1}|}-\frac{1}{|V_{w,2}|}\bigg)(V_{w,1}\cdot\partial_{i}V_{w,2})V_{w,2}+\frac{(V_{w,1}-V_{w,2})\cdot\partial_{i}V_{w,2}}{|V_{w,2}|}V_{w,2}.$
@@
$\|\partial_{i}(\mathcal{T}_{w,1}-\mathcal{T}_{w,2})\|_{L^{2}}\lesssim(\|V_{w,1}\|_{L^{\infty}}+\|V_{w,2}\|_{L^{\infty}})\|\partial_{i}\delta u\|_{L^{2}}+(\|\partial_{i}V_{w,1}\|_{L^{4}}+\|\partial_{i}V_{w,2}\|_{L^{4}})\|\delta u\|_{L^{4}}.$ (65)
Hence one has by using (64)
$|J_{8}|$ $\lesssim\lVert\nabla\delta u\rVert_{L^{2}}\bigg[\big(\lVert U_{w}-u_{1}\rVert_{L^{\infty}}+\lVert U_{w}-u_{2}\rVert_{L^{\infty}}\big)\lVert\nabla\delta u\rVert_{L^{2}}$ (66)
@@
Thus, we have derived the following from estimate (63),
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}).\end{gathered}$ (67)
Consequently, combining estimates (62) and (67) leads to
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}(1+\alpha^{2})\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}),\end{gathered}$ (68)
from which the stability (i.e. the continuous dependence on the initial data) and uniqueness of the strong solution follows. This concludes the proof of the global well-posedness of the intermediate model (13).
@@
We consider a sequence of solutions $\{(u_{\epsilon},\sigma_{\epsilon})\}$ of the intermediate system (13) such that $\epsilon\rightarrow 0$. Thanks to estimate (59), which is independent of $\epsilon$, by the Aubin-Lions lemma and Banach-Alaoglu theorem we have the following convergence results
$u_{\epsilon}$ $\overset{\ast}{\rightharpoonup}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (69)
$\partial_{t}u_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\partial_{t}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (70)
$u_{\epsilon}$ $\rightarrow u$ strongly in $~C([0,T];H^{1}(\mathbb{T}^{2})),$ (71)
$\sigma_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (72)
$\sigma_{\epsilon}$ $\rightarrow\sigma$ strongly in $~C([0,T];H^{2}(\mathbb{T}^{2})),$ (73)
$\partial_{t}\sigma_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\partial_{t}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (74)
where the limits $u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\partial_{t}u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and $\partial_{t}\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$ satisfy estimate (59). It is left to verify that $(u,\sigma)$ is a solution to the Voigt-EVP model (12).
@@
where we have used Lemma 2.3, the convergence $u_{\epsilon}\rightarrow u$ in $C([0,T];H^{1}(\mathbb{T}^{2}))$ and the elementary inequality $\sqrt{y+z}\leq\sqrt{y}+\sqrt{z}$ for $y,z\geq 0$. It is clear that the linear terms in equations (13a)-(13b) converge. We now show that the nonlinear terms in the equation for the stress tensor converge. In particular, for all $t\in[0,T]$ we have
@@
In a similar fashion, one can show that the terms $\frac{3}{2}\frac{\sqrt{\lvert D(u_{\epsilon})\rvert^{2}+\epsilon^{2}}}{P}\operatorname{Tr}\sigma_{\epsilon}\mathbb{I}_{2}$ and $\frac{\sqrt{\lvert D(u_{\epsilon})\rvert^{2}+\epsilon^{2}}}{2}\mathbb{I}_{2}$ converge to $\frac{3}{2}\frac{\lvert D(u)\rvert}{P}\operatorname{Tr}\sigma\mathbb{I}_{2}$ and $\frac{\lvert D(u)\rvert}{2}\mathbb{I}_{2}$ in $L^{\infty}((0,T);L^{2}(\mathbb{T}^{2}))$. Finally, we show that the ocean stress term $\mathcal{T}_{w}^{\epsilon}$ converges. We find
@@
$\lesssim\;$ $\lVert u-u_{\epsilon}\rVert_{L^{4}}\xrightarrow[]{\epsilon\rightarrow 0}0,$
@@
To show the (global) well-posedness, one can repeat similar arguments as in Section 3.3. If the Voigt-EVP model would have two strong solutions $(u_{1},\sigma_{1}),(u_{2},\sigma_{2})\in C([0,T];H^{2}(\mathbb{T}^{2}))\times C([0,T];H^{3}(\mathbb{T}^{2}))$, then we again consider the difference
Similarly to before, one can check that the difference $(\delta u,\delta\sigma)$ satisfies the following system
$\partial_{t}\delta u=\nabla\cdot\delta\sigma+\mathcal{T}_{w,1}-\mathcal{T}_{w,2}+\Omega(\delta u)^{\perp},$ (75a)
$\frac{1}{\mathcal{E}}\partial_{t}\big(\delta\sigma-\alpha^{2}\Delta\delta\sigma\big)+\frac{4\mathcal{D}_{1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}$ $\mathbb{I}_{2})$ (75b)
$+\frac{\mathcal{D}_{1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}+\frac{\mathcal{D}_{1}-\mathcal{D}_{2}}{2}$ $\mathbb{I}_{2}=D(\delta u),$ (75c)
where we have introduced the following notation for $i=1,2$
$\mathcal{T}_{w,i}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{i}\rvert\big[(U_{w}-u_{i})\cos\theta+(U_{w}-u_{i})^{\perp}\sin\theta\big],$ (76)
$\mathcal{D}_{i}$ $=\lvert D(u_{i})\rvert$. (77)
Subsequently, by proceeding entirely analogously to section 3.3 (the only difference being that we use Lemma 2.3 in the case $\epsilon=0$ rather than $\epsilon>0$), we can obtain the analogous version of estimate (68), which is given by
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}(1+\alpha^{2})\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}).\end{gathered}$ (78)
From this estimate we can deduce the stability and uniqueness for the Voigt-EVP model. This finishes the proof of Theorem 1.2.
@@
We remark that the proof of the convergence of the solutions in the viscosity cutoff limit $\epsilon\rightarrow 0$, is in some sense related to the approach in Ref. [3], in which the local well-posedness of a model for non-Newtonian incompressible flows with power-law stress tensor was established uniformly in the cutoff parameter of the viscosity.
@@ 5 Discussion and Conclusion
In this paper we commenced the rigorous analytical study of one of the fundamental models of sea-ice dynamics, namely the elastic-viscous-plastic (EVP) sea-ice model. We introduced a regularised version of the EVP model, by means of the Voigt regularisation, which is inspired by the original elastic regularisation of the Hibler model in Ref. [39] (as both regularisations formally do not impact the asymptotic dynamics). We have shown this Voigt-EVP model to be globally well-posed. One interesting question, which we leave to future work, is whether the Voigt-EVP model also has a suitable notion of weak solution.
Due to the (formal) linear ill-posedness of the EVP model as discussed in Section 1.2, it seems necessary to introduce a regularised version of the EVP model of the type (12a)-(12c), in order to obtain a well-posed sea-ice model. Finally, in order to conclude the proof of the global well-posedness we took the limit of the viscosity cutoff parameter $\epsilon$ going to zero (while keeping the Voigt regularisation), and obtain that the Voigt-EVP model is globally well-posed.
Moreover, we would like to mention that over time several related models to the EVP model have been developed in the sea-ice literature, see for example Refs. [7, 38]. The approach developed here will also have potential applications in the rigorous study of several of these related sea-ice models.
@@
If one includes an advection term in the momentum equation of the intermediate system (13a)-(13c), by applying the same methods as in this paper, one can prove the local well-posedness of the intermediate system by using similar estimates as in this paper. Note however that the existence time will most likely depend on $\epsilon$ (unlike for Corollary 1.4). This therefore prohibits taking the limit $\epsilon\rightarrow 0$, as we have done in the proof of Theorem 1.2.
@@
We remark that the replacement of $\overline{\mathcal{D}}$ by $\lvert D(u)\rvert$ in the Voigt-EVP model is done merely for the sake of convenience. In fact, by using the approach from this paper one can prove the global well-posedness of the model (13a)-(13c) in which $\mathcal{D}_{\epsilon}$ has been replaced by
@@
$\left\lVert\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right\rVert_{L^{2}}\lesssim\lVert\nabla u\rVert_{L^{2}}+\epsilon,\quad\left\lVert\nabla\left(\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right)\right\rVert_{L^{2}}\lesssim\lVert\Delta u\rVert_{L^{2}}.$ (79)
@@
We also note that, in addition to (14), different regularisations of the strain rate have been used in the sea-ice literature. For example, one can set maximal values for the shear and bulk viscosities, as is done in Ref. [36]. This means that, instead of $\mathcal{D}_{\epsilon}$, one uses
$\max\{\overline{\mathcal{D}},\epsilon\}=\frac{\overline{\mathcal{D}}+\epsilon}{2}+\frac{1}{2}\lvert\overline{\mathcal{D}}-\epsilon\rvert$. (80)
@@
$\overline{\mathcal{D}}_{\epsilon,\gamma}\coloneqq\frac{\overline{\mathcal{D}}+\epsilon+\gamma}{2}+\frac{1}{2}\sqrt{\lvert\overline{\mathcal{D}}-\epsilon\rvert^{2}+\gamma^{2}}.$ (81)
@@
It is straightforward to repeat the proof of Theorem 1.2 to obtain a global strong solution of the Voigt-EVP model with the viscosity regularisation from Eq. (80) by taking the limit $\gamma\rightarrow 0$. Moreover, one can also choose to regularise the viscosities by means of the hyperbolic tangent, as was introduced in Ref. [52], see also Ref. [18], but again one can treat this regularisation by similar methods as in this paper. Generally speaking, one can treat any $C^{2}$ approximation of $\overline{\mathcal{D}}$ by relying on the approach of this paper.